Police response time to an emergency call is the difference between the time the call is first received by the dispatcher and the time a patrol car radios that it has arrived at the scene. Over a long period of time, it has been determined that the police response time has a normal distribution with a mean of 9.6 minutes and a standard deviation of 2.3 minutes. For a randomly received emergency call, find the following probabilities. (Round your answers to four decimal places.)
(a) the response time is between 5 and 11 minutes


(b) the response time is less than 5 minutes


(c) the response time is more than 11 minutes
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 9.6
std deviation   =σ= 2.300

a)

probability =P(5<X<11)=P((5-9.6)/2.3)<Z<(11-9.6)/2.3)=P(-2<Z<0.61)=0.7291-0.0228=0.7063

b_)

probability =P(X<5)=(Z<(5-9.6)/2.3)=P(Z<-2)=0.0228

c)

probability =P(X>11)=P(Z>(11-9.6)/2.3)=P(Z>0.61)=1-P(Z<0.61)=1-0.7291=0.2709